# Not (yet) the whole story: Evaluating Visual Storytelling Requires More than Measuring Coherence, Grounding, and Repetition

### v2

@@ 1 Introduction
Nevertheless, we argue that measuring the degree of coherence or visual grounding of a story may not be sufficiently informative, as there are no standard conventions that determine the preferable level of such properties. To address this issue, in this work, we first propose an evaluation method that assesses the quality of generated stories in terms of their distance from human-written stories along several relevant dimensions, each measured by an available reference-free metric. Using this method, we evaluate a range of models on the visual storytelling task, including models specifically designed and trained for this task, as well as—for the first time—foundation models pre-trained to achieve general-purpose language and vision abilities, which we test in a zero-shot manner. We show that LLaVA (Liu et al., 2024), a powerful foundation model, performs best on the task, but only slightly so than TAPM (Yu et al., 2021), a model designed for visual storytelling which is 50 times smaller than LLaVA. Second, given insights derived from our proposed distance-based evaluation method, we upgrade the visual and language components on TAPM, resulting in a model that achieves comparable performance to LLaVA with a significantly lower number of parameters.
@@
  • We use our evaluation to assess the stories generated by various visual storytelling-specific and foundation models; we find that a foundation model, LLaVA, achieves the best result (lowest distance).
@@ 4.4 Results
Regarding the two best performing models, LLaVA and TAPM, two points are worth highlighting. First, despite a huge difference in model size—LLaVA is a powerful 7.5B parameter foundation model, while TAPM is 50 times smaller—TAPM’s $\text{d}_{HM}$ value is only slightly higher than LLaVA’s. Second, LLaVA outperforms TAPM with respect to coherence and visual grounding. We hypothesize that this advantage is due to LLaVA’s more powerful language and vision backbone models. In the next section, we leverage the modular architecture of TAPM to test this hypothesis.
@@
AREL the friends were having a great time at the party. we had a lot of fun. we had a lot of beer. we all had a great time and had a great time. we had a great time.
@@
BLIP-2 a group of people standing outside drinking beer and talking on cell phones . . a hot dog is being cooked on a stick over a fire . . a group of people standing in a room with a woman making a surprised face
Figure 2: An example from the VIST test set with corresponding model-generated stories.
@@ 5.4 Results
Figure 3 shows the $\text{d}_{HM}$ values for the LLaVA model and for all the different versions of the TAPM model on both the VIST and the VWP datasets. Firstly, we observe that compared to the original TAPM model, (+Llama 2) generates stories that are closest to human stories in terms of both coherence and non-redundancy, whereas stories by the (+ViT) version are closest along the dimension of visual grounding. These results are in line with our intuitions regarding the influence that both modalities have on different aspects of visual story evaluation. For this analysis, we also considered a version of the TAPM model in which both the language and vision components are jointly updated (+Llama 2, +ViT). However, (+Llama 2, +ViT) consistently under-performed on our distance measure $\text{d}_{HM}$ compared to the (+Llama 2) and (+ViT) versions, and was only marginally better than the original TAPM model. Despite the significant difference in the number of parameters, we notice that (+Llama 2) achieves performance on par with LLaVA (see Figure 4).
We observe similar results for the TAPM models on the VWP dataset—(+ViT) obtains the lowest distance in terms of visual grounding and (+Llama 2) obtains the lowest distance in terms of both coherence and non-redundancy. Furthermore, on the VWP dataset, the (+Llama 2) model achieves the overall lowest $\text{d}_{HM}$, performing better than the LLaVA model. These results quantitatively indicate that the stories generated by the models are close to the corresponding human-written stories. To better understand if the metrics are capable of effectively comparing stories generated by models with human-written ones, we conduct a qualitative evaluation, that we describe in the next section.
@@ 6 Qualitative Analysis and Discussion
Yet, preference for human-written stories suggests that a good story involves more than just visual grounding, coherence, and limited repetition. To get more insight into this issue, we asked the annotators in the study to describe the properties they considered when comparing stories. Some annotators reported that they employed their subjective world knowledge in differentiating between creative and hallucinated stories. Also, stories that describe events without an overarching narrative were judged as bad despite being locally coherent, well-grounded, and non-repetitive. In addition, annotators generally preferred stories that contained phrases or sentences expressing emotions. We provide examples of these cases in Appendix C.
@@ Appendix C Human Evaluation
Figure 9: Example depicting emotions: all annotators evaluated S2 as better than S1.
@@
Figure 10: Example with hallucinations: 4 annotators (out of 5) either selected S1 as better or evaluated S2 as bad.
@@
Figure 11: Example with an overarching narrative: all annotators evaluated S1 as better than S2.

### v3

@@ 1 Introduction
Nevertheless, we argue that measuring the degree of coherence or visual grounding of a story may not be sufficiently informative, as there are no standard conventions that determine the preferable level of such properties. To address this issue, in this work, we first propose an evaluation method that assesses the quality of generated stories in terms of their distance from human-written stories along several relevant dimensions, each measured by an available reference-free metric. Using this method, we evaluate a range of models on the visual storytelling task, including models specifically designed and trained for this task, as well as—for the first time—foundation models pre-trained to achieve general-purpose language and vision abilities, which we test in a zero-shot manner. We show that LLaVA (Liu et al., 2024), a powerful foundation model, performs best on the task, but only slightly better than TAPM (Yu et al., 2021), a model designed for visual storytelling which is 50 times smaller than LLaVA. Second, given insights derived from our proposed distance-based evaluation method, we upgrade the visual and language components on TAPM, resulting in a model that achieves comparable performance to LLaVA with a significantly lower number of parameters.
@@
  • We use our evaluation to assess the stories generated by various visual storytelling-specific and, for the first time in the community, general-purpose foundation models; we report the novel finding that a foundation model, LLaVA, achieves the best result (lowest distance) when prompted under a novel setting.
@@ 4.4 Results
Regarding the two best-performing models, LLaVA and TAPM, two points are worth highlighting. First, despite a huge difference in model size—LLaVA is a powerful 7.5B parameter foundation model, while TAPM is 50 times smaller—TAPM’s $\text{d}_{HM}$ value is only slightly higher than LLaVA’s. Second, LLaVA outperforms TAPM with respect to coherence and visual grounding. We hypothesize that this advantage is due to LLaVA’s more powerful language and vision backbone models. In the next section, we leverage the modular architecture of TAPM to test this hypothesis.
@@
C=0.993, G=0.933, R=0.968
AREL the friends were having a great time at the party. the fire was <UNK> and <UNK>. the fire was <UNK> and <UNK>. the guys were having a great time. we all had a great time and had a great time.
C=0.562, G=0.348, R=0.670
@@
C=0.999, G=0.574, R=0.841
BLIP-2 a group of people standing outside drinking beer and talking on cell phones . . a hot dog is being cooked on a stick over a fire . . a group of people standing in a room with a woman making a surprised face.
C=0.294, G=1.024, R=0.884
Figure 2: An example from the VIST test set with corresponding model-generated stories and scores from individual evaluation metrics (Coherence: C, Visual grounding: G, Repetition: R).
@@ 5.4 Results
Figure 3 shows the $\text{d}_{HM}$ values for the LLaVA model and for all the different versions of the TAPM model on both the VIST and the VWP datasets. Firstly, we observe that compared to the original TAPM model, (+Llama 2) generates stories that are closest to human stories in terms of both coherence and repetition, whereas stories by the (+ViT) version are closest along the dimension of visual grounding. These results are in line with our intuitions regarding the influence that both modalities have on different aspects of visual story evaluation. For this analysis, we also considered a version of the TAPM model in which both the language and vision components are jointly updated (+Llama 2, +ViT). However, (+Llama 2, +ViT) consistently under-performed on our distance measure $\text{d}_{HM}$ compared to the (+Llama 2) and (+ViT) versions, and was only marginally better than the original TAPM model. Despite the significant difference in the number of parameters, we notice that (+Llama 2) achieves performance on par with LLaVA (see Figure 4).
We observe similar results for the TAPM models on the VWP dataset—(+ViT) obtains the lowest distance in terms of visual grounding and (+Llama 2) obtains the lowest distance in terms of both coherence and repetition. Furthermore, on the VWP dataset, the (+Llama 2) model achieves the overall lowest $\text{d}_{HM}$, performing better than the LLaVA model. These results quantitatively indicate that the stories generated by the models are close to the corresponding human-written stories. To better understand if the metrics are capable of effectively comparing stories generated by models with human-written ones, we conduct a qualitative evaluation, that we describe in the next section.
@@ 6 Qualitative Analysis and Discussion
Yet, preference for human-written stories suggests that a good story involves more than just visual grounding, coherence, and limited repetition. To get more insight into this issue, we asked the annotators in the study to describe the properties they considered when comparing stories. Some annotators reported that they employed their subjective world knowledge in differentiating between creative and hallucinated stories. Also, stories that describe events without an overarching narrative were judged as bad despite being locally coherent, well-grounded, and non-repetitive. In addition, annotators generally preferred stories that contained phrases or sentences expressing emotions. We provide examples of these cases in Appendix C, which illustrate the limitations of current evaluation metrics in fully capturing hallucinations, relevant emotions, and creative expressions in the generated stories. Therefore, there is scope for a lot of further work in this evaluation domain.
@@ Appendix C Human Evaluation
C=0.504, G=0.815, R=0.901
Figure 9: Example depicting emotions: all annotators evaluated story S2 as better than story S1. C, G, and R denote the coherence, visual grounding, and repetition scores respectively.
@@
C=0.647, G=0.703, R=0.971
Figure 10: Example with hallucinations: 4 annotators (of 5) either selected story S1 as better or evaluated story S2 as bad. C, G, and R denote the coherence, visual grounding, and repetition scores respectively.
@@
C=0.996, G=1.576, R=0.867
Figure 11: Example with an overarching narrative: all annotators evaluated story S1 as better than story S2. C, G, and R denote the coherence, visual grounding, and repetition scores respectively.